14. Weights of a certain model of fully loaded gravel trucks follow a normal distribution with mean 

 = 6.4 tons and standard deviation  = 0.4 ton.

 

a) For a truck selected at random, what is the 

probability the weight will be 

less than 5.9 tons?

 

 

 

 

b) For a truck selected at random, what is the 

probability the weight will be 

more than 7 tons?

 

 

 

 

 

c) For a truck selected at random, what is the 

probability the weight will be 

between 5.4 and 6.4 tons?

 

 

 


1
Expert's answer
2021-05-24T15:35:12-0400

We have given that,

"\mu = 6.4"

"\sigma = 0.4"

a.) The probability the weight will be less than 5.9 tons

"=P(X<5.9)"


"=P(Z<\dfrac{5.9-6.4}{0.4})"


"=P(Z<-1.25) = 0.10"


b.) The probability the weight will be more than 7 tons

"=P(X>7)"


"=P(Z>\dfrac{7-6.4}{0.4})"


"=P(Z>1.5)"


"=0.10"


c.) The probability the weight will be between 5.4 and 6.4 tons

"P(5.4<X<6.4)"


"= P(\dfrac{5.4-6.4}{0.4}<Z<\dfrac{6.4-6.4}{0.4})"


"= P(-2.5<Z<0)"


"= 0.49"